OpenStudy (anonymous):

solve for x. 9(x-2)=5(2+x)

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

Distribute first: 9x-18=10+5x

OpenStudy (anonymous):

9(x-2)=5(2+x) left side: distribute the 9 to the terms inside the parenthesis right side: distribute the 5 to the terms inside the parenthesis

OpenStudy (anonymous):

i need help just give me the answer this is due tomorrow

OpenStudy (anonymous):

I won't give you a straight out answer, i will help you out and reach the answer tho :)

OpenStudy (anonymous):

9(x-2)=5(2+x) left side: distribute the 9 to the terms inside the parenthesis right side: distribute the 5 to the terms inside the parenthesis 9(x-2)=5(2+x) (9*x) - (9*2) = (5*2) + (5*x) 9x - 18 = 10 + 5x subtract 5x from both sides 9x - 18 = 10 + 5x -5x -5x --- ----- 4x - 18 = 10 add 18 to both sides 4x - 18 = 10 +18 +18 ----- ----- 4x = 28 divide by 4 on both sides 4x = 28 --- --- 4 4 x = ?

OpenStudy (anonymous):

If you do not understand a step, please do not hesitate and ask :)